# Some remarks on L-equivalence for cubic fourfolds and hyper-Kähler manifolds

### v1

@@ 4.2.2 Hilbert schemes of points on a K3 surface
Let $X$ be hyper-Kähler of K3[n] type, in [Taelman:derived_equivalences, beckmann2023derived] the authors define a Hodge lattice $\Lambda_{X}\subset\widetilde{H}(X,\mathbb{Q})$. As abstract lattices, $\Lambda_{X}\cong U\oplus H^{2}(X,\mathbb{Z})$ and its Hodge structure is such that $H^{2}(X,\mathbb{Z})^{\perp}\subseteq\Lambda_{X}^{1,1}$. If two manifolds of K3[n] type are D-equivalent, then there is a Hodge isometry $\Lambda_{X}\cong\Lambda_{X^{\prime}}$ by [beckmann2023derived, Theorem 9.2].
Lemma 4.15.
Let $X$ be a manifold of K3[n] type such that $\operatorname{End}(T(X))=\mathbb{Z}$ and $\operatorname{rk}T(X)\not=4$. Assume that $X^{\prime}$ is another manifold of K3[n] type L-equivalent to $X$ with $\operatorname{disc}T(X)=\operatorname{disc}T(X^{\prime})$. Assume that $D({\operatorname{NS}(X)})$ and $D({\operatorname{NS}(X^{\prime})})$ are isometric, then there is a Hodge isometry $\Lambda_{X}\cong\Lambda_{X^{\prime}}$.
@@
Since $\operatorname{disc}T(X)=\operatorname{disc}T(X^{\prime})$, $\operatorname{End}(T(X))=\mathbb{Z}$ and $\operatorname{rk}T(X)\not=4$, then by 2.7 and 2.9 we have a Hodge isometry $T(X)\cong T(X^{\prime})$. From the fact that $U=H^{2}(X,\mathbb{Z})^{\perp}\subseteq\Lambda_{X}^{1,1}$ we have $\operatorname{rk}\Lambda_{X}^{1,1}\geq l(\Lambda_{X}^{1,1})+2$. Moreover, since $\operatorname{NS}(X)$ and $\operatorname{NS}(X^{\prime})$ have the same genus then also $\Lambda_{X}^{1,1}$ and $\Lambda_{X^{\prime}}^{1,1}$ have the same genus. We are then in the place to apply 2.5 to get a Hodge isometry $\Lambda_{X}\cong\Lambda_{X^{\prime}}$.
Moreover, in some specific cases there is a better description of Hodge isometries induced by derived equivalence.
Proposition 4.16.
Let $X=S^{[n]}$ for n1n-1 a prime power and a K3 surface $S$ with $U\subseteq\operatorname{NS}(S)$, and assume that $\operatorname{End}(T(X))=\mathbb{Z}$ and $\operatorname{rk}T(X)\not=4$. If $X^{\prime}$ is a manifold of K3[n] type which is L-equivalent to $X$ and $\operatorname{disc}T(X)=\operatorname{disc}T(X^{\prime})=m$ is coprime to 2(n1)2(n-1), then $X$ and $X^{\prime}$ are D-equivalent.
Proof.
In this case we know from [beckmann2023derived, Theorem 9.8] that any Hodge isometry $\Lambda_{X}\cong\Lambda_{X}$ preserving the orientation of a positive-definite real $4$-plane is induced by a derived equivalence $D^{b}(X)\cong D^{b}(X^{\prime})$. It suffices then to show that 4.15 can be applied in this situation giving a Hodge isometry $\Lambda_{X}\cong\Lambda_{X^{\prime}}$ and that this Hodge isometry can be chosen to be orientation-preserving.
Firstly, we need to show that $D({\operatorname{NS}(X)})$ is isometric to $D({\operatorname{NS}(X^{\prime})})$. We apply 2.7 to obtain a Hodge isometry $T(X)\cong T(X^{\prime})$. By hypothesis, the discriminant group $D({T(X^{\prime})})$ does not contain any subgroup isomorphic to $\mathbb{Z}/2(n-1)\mathbb{Z}$ so that the primitive embedding $T(X^{\prime})\subset H^{2}(X^{\prime},\mathbb{Z})$ has trivial embedding subgroup and then the discriminant of the orthogonal complement $\operatorname{NS}(X^{\prime})$ is given by $D({\operatorname{NS}(X^{\prime})})\cong D({T(X^{\prime})})\oplus D({[-2(n-1)]})$, which is isometric to the discriminant form of $\operatorname{NS}(X)$. We are now in the hypothesis of 4.15, which gives an Hodge isometry $\Lambda_{X}\cong\Lambda_{X^{\prime}}$.
We finally observe that the Hodge isometry can be chosen to be orientation preserving. In fact, from $U\subseteq\operatorname{NS}(X)$ we have that $\operatorname{rk}\Lambda_{X}^{1,1}\geq l(D({\Lambda_{X}^{1,1}}))+4$ so that by [Nikulin_integral, Corollary 1.13.5] the lattice $\Lambda_{X}^{1,1}$ splits a copy of $U$. Hence, if the Hodge isometry $\Lambda_{X}\cong\Lambda_{X^{\prime}}$ is reversing the orientation, after composing with the isometry acting as $\mathrm{id}$ on the copy of $U$ and trivially on the complement, we obtain a Hodge isometry preserving the orientation. ∎

### v2

@@ 4.2.2 Hilbert schemes of points on a K3 surface
We now consider manifolds of K3[n] type that are birational so Hilbert schemes of points on a K3 surface.
Proposition 4.14.
Let $X$ birational to $S^{[n]}$ for a K3 surface $S$ with $U\subseteq\operatorname{NS}(S)$, and assume that $\operatorname{End}(T(X))=\mathbb{Z}$ and $\operatorname{rk}T(X)\not=4$. If $X^{\prime}$ is a manifold of K3[n] type which is L-equivalent to $X$ and $\operatorname{disc}T(X)=\operatorname{disc}T(X^{\prime})=m$ is coprime to 2(n1)2(n-1), then $X$ and $X^{\prime}$ are D-equivalent.
@@
Since $\operatorname{disc}T(X)=\operatorname{disc}T(X^{\prime})$, $\operatorname{End}(T(X))=\mathbb{Z}$ and $\operatorname{rk}T(X)\not=4$, then by 2.7 and 2.9 we have a Hodge isometry $T(X)\cong T(X^{\prime})$.
Firstly, we want to show that $D({\operatorname{NS}(X)})$ is isometric to $D({\operatorname{NS}(X^{\prime})})$. By hypothesis, the discriminant group $D({T(X^{\prime})})$ does not contain any subgroup isomorphic to $\mathbb{Z}/2(n-1)\mathbb{Z}$ so that the primitive embedding $T(X^{\prime})\subset H^{2}(X^{\prime},\mathbb{Z})$ has trivial embedding subgroup and then the discriminant of the orthogonal complement $\operatorname{NS}(X^{\prime})$ is given by $D({\operatorname{NS}(X^{\prime})})\cong D({T(X^{\prime})})(-1)\oplus D({[-2(n-1)]})$, which is isometric to the discriminant form of $\operatorname{NS}(X)\cong\operatorname{NS}(S)\oplus[-2(n-1)]$ from the fact that $D(\operatorname{NS}(S))\cong D(T(S))(-1)\cong D(T(X))(-1)$.
From the fact that $U\subseteq\operatorname{NS}(X)$ we have $\operatorname{rk}\operatorname{NS}(X)\geq l(\operatorname{NS}(X))+2$. Moreover, since $\operatorname{NS}(X)$ and $\operatorname{NS}(X^{\prime})$ have the same genus we are in the place to apply 2.5 to get a Hodge isometry $H^{2}(X,\mathbb{Z})\cong H^{2}(X^{\prime},\mathbb{Z})$. Up to composing with $\mathrm{id}$ the Hodge isometry can be chosen to be orientation-preserving. Moreover, since $\operatorname{End}(T(X))=\mathbb{Z}$ and $O(\operatorname{NS}(X))\to O(D(\operatorname{NS}(X)))$ is surjective with the fact that $D({\operatorname{NS}(X)})\cong D({T(X)})(-1)\oplus D({[-2(n-1)]})$, the Hodge isometry can be chosen to act as $\pm 1$ on the discriminant and hence by [markman2011survey, Corollary 9.9] we have that $X$ and $X^{\prime}$ are birational. Now, by [maulik2025d, Theorem 1.2] we know that they are also D-equivalent.